### v1

# Uncovering LLM-Generated Code: A Zero-Shot Synthetic Code Detector via Code Rewriting

Abstract.
Large Language Models (LLMs) have exhibited remarkable proficiency in generating code. However, the misuse of LLM-generated (Synthetic) code has prompted concerns within both educational and industrial domains, highlighting the imperative need for the development of synthetic code detectors. Existing methods for detecting LLM-generated content are primarily tailored for general text and often struggle with code content due to the distinct grammatical structure of programming languages and massive ”low-entropy” tokens. Building upon this, our work proposes a novel zero-shot synthetic code detector based on the similarity between the code and its rewritten variants. Our method relies on the intuition that the differences between the LLM-rewritten and original codes tend to be smaller when the original code is synthetic. We utilize self-supervised contrastive learning to train a code similarity model and assess our approach on two synthetic code detection benchmarks. Our results demonstrate a notable enhancement over existing synthetic content detectors designed for general texts, with an improvement of 20.5% in the APPS benchmark and 29.1% in the MBPP benchmark.
Synthetic Code Detector, Code Rewriting, Code Large Language Models
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2024; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
1. Introduction
Large Language models (LLM) pretrained on code have shown remarkable capability in understanding and generating code (Chen et al., 2021; Fried et al., 2022; Li et al., 2023; Nijkamp et al., 2022). Those Code LLMs can function as professional coding assistants for programmers, offering intelligent code completion and document generation capabilities, such as Github Copilot. Furthermore, recent advancements in instruction tuning and AI alignment research (Wei et al., 2022a; Ouyang et al., 2022) have facilitated the development of general-purpose conversational LLMs (Chowdhery et al., 2022; OpenAi, 2022; OpenAI, 2023; Touvron et al., 2023). GPT-4, one of the most impressive conversational LLM released by OpenAI, can provide high-quality responses to general human requests in a zero-shot manner, including requests for generating code implementation for detailed coding specifications (Bubeck et al., 2023).
The breakthrough of LLM has dramatically improved the efficiency of the coding process and significantly lowered the barrier to entry for programming. However, it also raises concerns about the misuse of LLM-generated code. In programming education, students have already used LLMs to write solutions for coding assignments and exams (Kazemitabaar et al., 2023; Denny et al., 2024). A recent study has indicated that GPT-4 can achieve the average human level performance in solving LeetCode problems (Bubeck et al., 2023), which makes cheating with the use of LLMs in programming exams very attractive for students, and also makes it difficult to assess students’ programming abilities in education. Besides educational domain, in real industrial applications, there is a high demand for code security, and LLM-generated code often contains security vulnerabilities (He and Vechev, 2023). An evaluation (Pearce et al., 2022) discovered that, in various security-relevant scenarios, 40% of Copilot-generated programs contain dangerous vulnerabilities. Therefore, if the code is LLM-generated, it requires even more rigorous review.
Figure 1. Token Entropy Distribution of Natural Texts and Code. We use 500 documents from XSum (Narayan et al., 2018) dataset for token entropy estimation and 770 human-written code for code token entropy estimation.
Building upon the aforementioned critical real-world scenarios, developing a synthetic code detector is curcial for addressing concerns regarding the misuse of LLM-generated code. Although there have been some efforts to detect LLM-generated pure text (Gehrmann et al., 2019; Mitchell et al., 2023; Zellers et al., 2019; Ippolito et al., 2020; Zhong et al., 2020; Pu et al., 2023), none of them have specifically focused on code content. The question of whether existing detection methods for general texts remain effective when applied to code content is still unknown. However, unfortunately, the answer to this question is negative. Even state-of-the-art detection methods, such as GLTR (Gehrmann et al., 2019) and DetectGPT (Mitchell et al., 2023), exhibit a significant decline in detection performance when applied to the code domain, as revealed by our experiments. Our findings also align with a recent evaluation study of AIGC detectors on code content (Wang et al., 2023).
Figure 2. Large Language Models can only make minor modifications to the synthetic code. We use the StarChat model (Li et al., 2023) to rewrite the original code and observe that the similarity between the rewritten and original code is much higher when the original code is sampled from GPT-3.5-Turbo.
Through our meticulous experiments, we identify the primary reason lies in the core logic of existing state-of-the-art methods for text detection, which is based on the log probability of tokens. Large Languages models assign higher log probability to the tokens they generate because these tokens are considered by the model to the most confident in its self-generated content. Consequently, existing detection methods leverage this characteristic to differentiate between model-generated text and human-written text, achieving high detection accuracy. However, in the domain of code, this approach exposes its inherent weaknesses. The main reason is that, compared to natural language, code exhibits a more uniform grammatical structure. In specific programming language and the given context, the subsequent code token is often deterministic or has limited options. For example, in the Python language, functions always start with ”def”, and a ”:” must follow ”def funciton_name()”. We refer to these tokens as ”low-entropy” tokens. However, there exists a large number of such ”low-entropy” tokens in code, reducing the actual effective tokens for SOTA text detection methods to differentiate between human-written code and model-generated code, resulting a statistically insignificant differences in distinguishing between these two types of code. We investigate the code token extropy distribution and the token entropy distribution (both estimated by Llama-13B) in Figure 1. The distrubution of text token entropy is more dispersed than code, and more than 60% of code tokens have entropy smaller than 1.
Therefore, to address the detection challenge between model-generated code and human-written code in the code domain, it is essential to abandon the current token log probabilities perspective. Adopting a more holistic, global view, on the other hand, opens up our thinking. Through an essential observation, as illustrated in Figure 2, that Large language Model, when tasked with rewriting model-generated code produced by either themselves or other LLMs, tend to exhibit remarkably similar coding patterns in their rewritten output compared to the original code. Conversely, when tasked with rewriting human-written code, the LLM-rewritten code diverges more significantly from the original code. In Figure 2, in the top left corner is the original code (model-generated), which, after beging rewritten by LLMs, yields the code in the top right corner. It’s evident that the two are strikingly similar, with matching portions highlighted in yellow. Conversely, the bottom left corner showcases a human-written code. to the same LLM rewriting process, we obtain the code in the bottom right corner, where we observe a substantial number of differences.
This observation inspired our hypothesis: the similarity between the original and rewritten code can be used as an indicator for detecting synthetic code. Furthermore, the generation of rewritten code and the utilization of similarity as an indicator allows detection methods to be independent of likelihood estimation, thereby addressing the root cause of failure in inaccurate likelihood estimation for code content. Based on this, we propose a zero-shot model-generated code detector using Code Rewriting and similarity Measurement.
Our approach comprises two crucial steps: Code Rewriting and Similarity Measurement. When confronted with a piece of code, whether model-generated or human-written, we first employ a large language model to rewrite it using an appropriate prompt template. This process yields both the original code and its corresponding rewritten version. Subsequently, to accurately gauge the similarity between the original and rewritten code, we develop a code similarity model that predicts a similarity score for the given code pair. Finally, employing the cosine similarity metric, we estimate the expected similarity score by sampling m rewritten codes following the identical procedure and averaging their scores.
We constructed two model-generated code detection benchmarks utilizing APPS (Hendrycks et al., 2021) and MBPP (Austin et al., 2021), representing use cases in the framework of synthetic code exploitation. To confirm that our intuition and hypothesis are universally applicable to all types of code generation tools, we generate synthetic code using two open-sourced LLMs (CodeLlama (Rozière et al., 2023), StarChat (Li et al., 2023)) and two proprietary code generation APIs (GPT-3.5, GPT-4) for detection.The experimental results demonstrate that the methods proposed in our study notably outperform SOTA benchmarks. Specially, our approach achieve 20.5% improvement in detection performance on the APPS dataset and 29.1% improvement in MBPP dataset. Furthermore, from a practicality and minimal resource requirement standpoint, our method only necessitates the capability to perform inference with or access to their APIs, in contrast to the previous text detection methods that require knowledge of token log probabilities.
To facilitate further research, we have made our dataset, code and trained code similarity model checkpoint publicly available. Overall, the main contributions of this paper can be outlined as follows:
  • We have identified a performance gap when applying synthetic content detectors designed for general texts to code content. Following a detailed analysis, we have uncovered the reasons for this discrepancy.
  • We proposed a zero-shot synthetic content detection tool specially designed for code content, applicable to both open-sourced code Large Language Models and closed-source LLMs like ChatGPT/GPT-4, which solely offer APIs.
  • We proposed a novel detection approach utilizing code rewriting and similarity measurement, which effectively addressed the unique challenges for synthetic code detection.
  • We conducted extensive experiments and analysis to demonstrate the effectiveness of our approach, showing significant improvements in accuracy compared to existing methods.
2. Related Work
2.1. LLMs for Code Generation
With the rapid rise in popularity of large language models, an increasing number of studies have begun focusing on code large language models aimed at achieving automatic software engineering. Those code large language models are continuely pre-trained on large-scale code corpus from an initial pre-trained model on text with auto denoising (Wang et al., 2021a) or causal language modeling task (Chen et al., 2021; Li et al., 2023; Nijkamp et al., 2022; Fried et al., 2022). A crucial capability of large code models is code generation. Code generation is a left-to-right decoding process that utilizes functional requirements or code context as a prompt. Pioneering works such as CodeX (Chen et al., 2021), AlphaCode (Li et al., 2022a), CodeGeex (Zheng et al., 2023), SantaCoder (Allal et al., 2023), and GPT-4 (Achiam et al., 2023) have all demonstrated powerful code generation capabilities. Recent advances in general purpose LLMs (Chowdhery et al., 2022; OpenAi, 2022; OpenAI, 2023; Touvron et al., 2023) further enhanced their understanding of human instructions, resulting in the generation of code that better meets the requirements and exhibits higher quality.
As code large language models continue to mature, more and more people from various industries are integrating these models into their daily lives. According to a community survey led by StackOverflow (2023), 44% of experienced developers and 55% of beginners already used AI coding assistant, and most of them use two tools, ChatGPT (83%) and Github Copilot (56%). Therefore, in certain areas such as education and industrial settings, automated detection of whether code is written by humans or generated by models has become increasingly important.
2.2. Detection of Synthetic Text
Detecting AI-generated (synthetic) text has been studied before the emergency of LLM. The main works on synthetic text detection have two lines. One research line formulates detection as a binary classification problem by collecting synthetic texts from generative models and training a supervised model based on pre-trained transformers (Zellers et al., 2019; Ippolito et al., 2020; Zhong et al., 2020) or other neural models (Bakhtin et al., 2019; Uchendu et al., 2020). Another research line detects synthetic text by designing zero-shot metrics. Those metrics measure the relationship between a given text and the text distribution of generative models. Gehrmann et al. (Gehrmann et al., 2019) claims that the synthetic texts are sampled from the head of generative models’ distribution, so the average log probability score under the generative model $\mathcal{G}$ of a given text, i.e., $\frac{1}{L}\sum_{i=1}^{L}log(P_{\mathcal{G}}(x_{i}|x_{1},...,x_{i-1}))$, can be a simple and effective zero-shot metric (called GLTR). Su et al. (2023) proposed two metric, LRR and NPR, where the former combines log-rand and log-likelihood to better magnify the differences between human-written and machine-written text, while the later is primarily based the idea that the log rank of machine-generated texts should be more sensitive to smaller perturbations. Mitchell et al. (2023) further improved the GLTR metrics by proposing that the synthetic texts lie at the probability curvature, i.e., the local maximum of the generative model’s distribution. They use the average token probability disparity between the given text and perturbed texts to detect whether the given text is located at the probability curvature of the generative model. DetectGPT4Code (Yang et al., 2023) introduces three key modifications to the original DetectGPT for adaptation to the code domain. Overall, all of these methods converge on the same idea: large language models tend to generate tokens with higher confidence. Therefore, effective differentiation can be achieved by statistically analyzing token log probabilities.
In addition to detection methods, some approaches have explored adding watermarks to the generated text during the generation process behind LLM services(Kirchenbauer et al., 2023; Yoo et al., 2023). Through watermark extraction, it becomes feasible to discern whether a given text originates from a model. However, it’s worth noting that the inclusion of a watermark algorithm could potentially compromise the quality of the generated text. Consequently, it cannot be presumed that all users or service providers will be inclined to produce watermarked text.
3. Method
3.1. Problem Definition
We focus on the zero-shot synthetic code detection problem. The synthetic code detection problem involves determining whether a code snippet $x$ is generated from a generative model or a API service (such as ChatGPT). The term ”zero-shot” refers to the scenario that we do not have access to a dataset with labeled synthetic code and human-written code for model training. All the previous zero-shot synthetic text detection methods, including GLTR and DetectGPT consider the ”white-box” setting, where log probability score of generative model is accessible. In the synthetic code detection problem, we argue that the ”white-box” assumption is too strong because in many cases, some commercial models only provide the generated content and not provide the log probability score. Therefore, we do not assume access to the log probability score and study the ”black-box” setting, where only generated content is available.
3.2. Overview
Figure 3. Prompt Template for Code Rewriting.
Our method is based on the intuition that every programmer has a unique coding style and tends to solve coding problems following their own routines and habits. Similarly, generative models can be likened to senior programmers, each with their distinctive coding style. This method is inspired by the code reviewing and refactoring process. In these stages, the original author of the code usually does not participate. Managers assign these tasks to other programmers to prevent the original authors from being stuck in their habitual routines, which might hinder them from making significant improvements. In a similar vein, Large Language Models tend to exhibit consistent code writing patterns due to inherent biases in their training data. Therefore, LLMs can be viewed as programmers with unique coding styles and habitual routines shaped by the common biases present in their training data.
Therefore, we hypothesize that when asking generative models to rewrite synthetic code, the differences between the rewritten and the original code tend to be smaller than when the original code is human written. Refer to Figure 2 for an intuitive understanding. Based on this hypothesis, we propose a zero-shot synthetic code detection method using Code Rewriting and Similarity Measurement. The overall design principle is: for a given code snippet $x$, we generate code rewriting multiple times using a generative LLM $\mathcal{G}$. Then we use a code similarity model $\mathcal{M}$ to measure the average similarity between rewritten code and original code $x$. We train a code similarity model $\mathcal{M}$ using self-supervised contrastive learning (Gao et al., 2021) on unlabeled code to measure the code similarity effectively. Consequently, we use the average similarity score as the detection metric. We summarize the proposed approach in Algorithm 1. Next, the two parts of Code Rewriting and Similarity Measurement will be explained in detail.
3.3. Code Rewriting
Given a code snippet $x$, we generate a rewriting of $x$ utilizing the chain of thought prompting method (Wei et al., 2022b). We prompt the generative model $\mathcal{G}$ with the original code snippet $x$ and instruct the model to initially generate an analysis and explanation of $x$, followed by a potential rewrite $x^{\prime}$. The intermediate code analysis can help the model to understand the original code and generate a valid rewrite according to the analysis. It should be noted that $x$ is normalized by omitting all comments and empty lines. Because comments are very easy to manipulate while maintaining their functionality. They can either be a source of noise or be utilized to evade detection.
To extract the code from the model’s response, we require the model to return the code in markdown format, and we subsequently remove the in-line comments during post-processing. The exact prompt template is depicted in Figure 3.
Input:
  $x$: code snippet;
  $\mathcal{G}$: generative model;
  $\mathcal{M}$: similarity model;
  $m$: number of rewriting;
  $\epsilon$: threshold;
Output:
  $true$: $x$ is generated by $\mathcal{G}$;
  $false$: $x$ is not generated by $\mathcal{G}$;
1
[…]
3 // generate m rewriting of x.
4 $score\leftarrow\frac{1}{m}\sum_{i}\frac{\mathcal{M}(x)\cdot\mathcal{M}(x^{\prime}_{i})^{T}}{\|\mathcal{M}(x)\|_{2}\times\|\mathcal{M}(x^{\prime}_{i})\|_{2}}$
5 // estimate the similarity score.
6 if $score>\epsilon$ then
7       return true   // x is generated by $\mathcal{G}$
8
9 end if
10else
11       return false   // x is not generated by $\mathcal{G}$
12 end if
Algorithm 1 Zero-shot Synthetic Code Detection
3.4. Similarity Measurement
To effectively measure the similarity between the rewritten code and the original code, we require a code similarity model $\mathcal{M}$, which can predict a similarity score $S$ for the input code pair. In the domain of AI for software engineering, code similarity learning is an activate area of research. Many studies address this problem by focusing on code representation learning. This involves generating dense semantic representations by leveraging various code structures (Feng et al., 2020; Ye et al., 2021; Wang et al., 2021b; Guo et al., 2021; Wang et al., 2022; Li et al., 2022c) and using vector distances to quantify similarity. We use GraphCodeBERT (Guo et al., 2021) as the similarity model since its model checkpoint is off-the-shelf, widely used and has been pre-trained on large-scale code.
To obtain better function-level representation from a pre-trained transformer, we employ self-supervised contrastive learning to enhance the sentence-level representation of GraphCodeBERT further. Following (Gao et al., 2021), we adapt the unsupervised SimCSE method to the code domain, using standard dropout as the data augmentation method for contrastive learning. For an input code snippet $x$, we use the last-layer hidden states of [CLS] token as code representation. During the unsupervised training stage, we add an MLP layer to obtain the final representation $\mathbf{h}$ of $x$.
Formally, for a batch of input code snippets $\{x_{i}\}_{i=1}^{N}$, we feed the batch to the GraphCodeBERT twice and get two embeddings, $\{\mathbf{h}_{i}\}_{i=1}^{N}$ and $\{\mathbf{h}^{\prime}_{i}\}_{i=1}^{N}$, with different dropout masks applied, the training objective of SimCSE is:
(1) $\mathcal{L}=-\frac{1}{N}\sum_{i=1}^{N}log\frac{e^{sim(\mathbf{h}_{i},\mathbf{h}^{\prime}_{i})/\tau}}{\sum_{j=1}^{N}e^{sim(\mathbf{h}_{i},\mathbf{h}^{\prime}_{j})/\tau}}$
Where $\tau$ is a temperature hyperparameter and is set to $0.1$. The $sim(\mathbf{h}_{i},\mathbf{h}^{\prime}_{i})$ is the cosine similarity $\frac{\mathbf{h}_{i}\mathbf{h}^{\prime}_{i}}{|\mathbf{h}_{i}||\mathbf{h}^{\prime}_{i}|}$.
The MLP layer is dropped after the unsupervised contrastive learning stage, and we only use the last-layer hidden states of the [CLS] token as snippet representation following (Gao et al., 2021). For a given code snippet $x$ and its rewriting $x^{\prime}$ sampled from $\mathcal{G}$ as introduced in Section 3.3, we obtain the final embedding of $x$ and $x^{\prime}$ by feeding them to the similarity model $\mathcal{M}$. It is worth noting that our framework is not dependent on a specific similarity model. $\mathcal{M}$ can be any other implementations that can accurately model the code similarity, such as OpenAI’s text embedding services (Greene et al., 2022).
We use the cosine similarity function and estimate the expectation of the similarity score by sampling $m$ rewritten code following the same procedure, as shown in Equation (2):
(2) $score(x)=\mathbb{E}_{x^{\prime}\sim\mathcal{G}(\cdot|x)}sim(\mathcal{M}(x),\mathcal{M}(x^{\prime}))$
A larger $m$ can result in a more accurate expectation estimation but at the cost of generating more rewritten code. Our experiments found that using only 4 rewrites can achieve outstanding detection performance.
4. Experiment Setting
4.1. Benchmarks
Considering there is no existing benchmark for evaluating synthetic code detectors. We build two synthetic code detection benchmarks in Python using APPS (Hendrycks et al., 2021) and MBPP (Austin et al., 2021). The choice of Python as the programming language for constructing our benchmark stems from several key considerations. Firstly, a significant portion of training data for prevalent code generation models like CodeGen (Nijkamp et al., 2022), Incoder (Fried et al., 2022), and StarCoder (Li et al., 2023) is constituted by Python code. These models exhibit notably superior code generation capabilities in Python than other languages. Consequently, the challenge and significance of detecting code generated by LLMs are most pronounced within the context of Python. Secondly, popular datasets used for code generation evaluation, such as HumanEval (Chen et al., 2021), MBPP (Austin et al., 2021), and APPS (Hendrycks et al., 2021), primarily provide Python test cases. We adopted Python as our chosen language to align with this prevailing practice. Furthermore, Python ranks among the most widely used programming languages, adding practical significance to validating our detection performance. Specifically, for the generation of synthetic code, we utilize the natural language description of each sample as a prompt and employ four widely-used code generation tools (CodeLlama (Rozière et al., 2023), StarChat (Tunstall et al., 2023), GPT-3.5 and GPT-4 (OpenAI, 2023)).
Choice of Generation Tools
We selected two open-sourced Code Large Language Models, CodeLlama-13B-Instruct (Rozière et al., 2023) and StarChat-Alpha (Tunstall et al., 2023), along with two proprietary generation APIs, GPT-3.5-turbo and GPT-4 (OpenAI, 2023). According to the survey (StackOverflow, 2023) conducted by StackOverflow, GPT-3.5-turbo is the most extensively used tool in daily programming. GPT-4 represents the state-of-the-art in code generation tools to date. Consequently, we have chosen these two models to typify proprietary generation APIs. CodeLlama (Rozière et al., 2023) and StarChat (Tunstall et al., 2023) are two potent open-sourced code generation tools, with their 34B versions exhibiting competitive performance when evaluated against GPT-3.5-turbo. We opted for the 13B and 15.5B versions due to resource constraints. Nonetheless, these versions also demonstrate superior generation performance.
Constructing the APPS benchmark.
APPS is a benchmark for code generation. Each sample in APPS is a coding problem taken from coding contest websites, e.g., CodeForces, LeetCode, and HackerRank, and contains problem descriptions and submitted solutions for the problem. The original APPS Benchmark comprises 5,000 test data and 5,000 training data instances. In our construction, we exclusively utilized the test data. The rationale for excluding the training data stems from concerns that the training data might have been incorporated into the training data of code LLMs. If the training data are using as prompts and those prompts are present in the LLM’s training data, it could potentially result in code generation that perfectly matches human-written code, leading to inaccuracies in evaluating the detection outcomes. Within the pool of 5,000 test data instances, we identified and removed prompts with web hyperlinks from the markdown format of coding challenges, as such links could affect code quality and are rarely used by developers in practice. This reduced the dataset to 3,846 instances. We then randomly selected 1,540 instances (40%) from this refined set and used 770 of them for synthetic code generation.
For those 770 problems, we generate synthetic code using the chain of thought prompting (Wei et al., 2022b) to improve the generation quality using four generation tools. Figure 4 shows the exact prompt used for code generation. The generation temperature is set to $0.7$, and $top\_p$ is set to 0.95. We removed all the comments and empty lines for the generated code. For the remaining 770 problems (1540 minus 770), we randomly sample a solution from all valid solutions as human-written code. The detailed benchmark statistics are listed in Table 1.
Generator CodeLlama StarChat GPT-3.5 GPT-4
# Sample 770/770 770/770 770/770 100/100
# Char 630/458 630/456 630/420 440/401
# Line 19.1/15.0 19.1/14.6 19.1/14.7 18.7/13.6
Table 1. APPS Benchmark Statistics. We count the averated number of chars and lines for both human-written/synthetic code for all versions of benchmark generated by four different code generation tools.
Constructing the MBPP benchmark.
MBPP is another Benchmark for evaluating AI code generation. It contains 1,000 crowd-sourced Python programming problems designed to be solvable by entry-level programmers, covering programming fundamentals, standard library functionality, etc. Compared to APPS, MBPP’s problems are more straightforward and frequently encountered in daily programming practice. We follow the same construction process as APPS. The detailed benchmark statistics are listed in Table 2.
Generator CodeLlama StarChat GPT-3.5 GPT-4
# Sample 233/233 233/233 233/233 100/100
# Char 256/192 256/251 256/265 267/238
# Line 10.0/6.6 10.0/8.2 10.0/9.2 10.4/8.4
Table 2. MBPP Benchmark Statistics.
All intermediate scripts used to construct the benchmark, as well as the final two benchmarks, are publicly available.
4.2. Similarity Model Training Details
We continue to train a code similarity model by exploiting unsupervised SimCSE from initial GraphCodeBERT as introduced in Section 3.4. In the SimCSE training stage, we collect 160k code snippets from CodeSearchNet (Husain et al., 2020) Python subset as training data. The GraphCodeBERT model trained by SimCSE is fixed for all subsequent experiments once training is complete. We use the Adam optimizer and set the maximum learning rate to $41e-$ 4 with linear decay. We use 4 NVIDIA 3090s with batch size 16 on each GPU and train the model for 5 epochs. In addition to GraphCodeBERT model, we also experimented with the UniXcoder (Guo et al., 2022) as another similarity model by following the same SimCSE training process.
4.3. Detector LLMs
Our approach necessitates a rewriting LLM, $\mathcal{G}$, to function as the detector LLM. Additionally, other zero-shot detectors for text, such as GLTR and DetectGPT, require a scoring LLM for likelihood estimation, as introduced in Section 2.2. We examine CodeLlama, StarChat-Alpha, and GPT-3.5 as detector LLMs for our method, while using CodeLlama and StarChat-Alpha for baseline comparison. We employ nucleus sampling for code rewriting, setting the top-$p$ to $0.95$ and the temperature to $0.8$.
4.4. Metrics
We utilize the Area Under the Receiver Operating Characteristic (AUROC) curve as the evaluation metric for all the detectors on our benchmarks, following DetectGPT (Mitchell et al., 2023). AUROC measures a classification model’s performance across various thresholds, providing a comprehensive evaluation of the method’s overall performance.
4.5. Baselines
We consider the following zero-shot detection methods using a surrogate model to approximate the true distribution of the generative model: $\mathbf{\log p(x)}$ (Gehrmann et al., 2019), LogRank, Rank, Entropy (Mitchell et al., 2023), LRR, NPR (Su et al., 2023) and DetectGPT (Mitchell et al., 2023). To adapt these detection metrics to code content and effectively approximate the true distribution of the generative code LLM, the surrogate score model should be replaced with LLMs trained on large-scale code content. In this study, we examine two open-sourced code LLMs: StarChat-Alpha and CodeLlama. We also select two supervised detectors and compare them to the zero-shot detectors: GPTZero (GPTZero, 2023) is a leading AI content detection servicen and is trained on millions of synthetic texts sampled from various generative models, including ChatGPT, GPT-4 and Bard. OpenAI-Detector (OpenAI, 2019), an open-sourced detector, is trained on texts sampled from GPT-2 and is based on Roberta-large (Liu et al., 2019).
$\mathbf{\log p(x)}$.
This method calculates the average log probability of each token. Code snippets with higher $\mathbf{\log p(x)}$ scores are more likely to be synthetic (Gehrmann et al., 2019).
LogRank and Rank.
These metrics use the averaged (log-) rank of tokens in the predicted distribution as a detection metric. Code snippets with lower ranks tend to be sampled from AI models.
Entropy.
We use the averaged predictive entropy of each token as another baseline following Mitchell et al. (2023). Synthetic code will be more ”in-distribution” for the generative model, leading to more confident predictions.
LRR and NPR
These two metrics, introduced by Su et al. (2023), are used in our study with the identical calculation formula as in the original paper without any modifications, ensuring fairness.
(3) $\text{LRR}=\left|\frac{\frac{1}{t}\sum_{i=1}^{t}\log p_{\theta}(x_{i}|x_{<i})}{\frac{1}{t}\sum_{i=1}^{t}\log r_{\theta}(x_{i}|x_{<i})}\right|$
where $r_{\theta}(x_{i}|x_{<i})\geq 1$ is the rank of token $x_{i}$ conditioned on the previous tokens.
(4) $\text{NPR}=\frac{\frac{1}{n}\sum_{p=1}^{n}\log r_{\theta}(\tilde{x}_{p})}{\log r_{\theta}(x)}$
where small perturbations are applied on the target text $x$ to produce the perturbed text $\tilde{x}_{p}$.
DetectGPT
DetectGPT (Mitchell et al., 2023) use the average token probability disparity between the given text and perturbed texts to detect whether the given text is located at the probability curvature of the generative model. To better adapt DetectGPT for the code domain, we made minor modification. Specifically, we replace the original perturbation model, T5-large (Raffel et al., 2020) with CodeT5-large (Wang et al., 2021a). All other settings are the same as in the original implementation.
GPTZero
It is a proprietary detection API (GPTZero, 2023). We request the provided API with the input code following their official documentation. We use the ”completely_generated_prob” field returned by their API as the detection score for AUROC calculation. ”completely_generated_prob” refers to the probability that the entire code snippet is generated by an AI model.
OpenAI-Detector
The finetuned Roberta-large model checkpoint provided by the authors is used for detection (OpenAI, 2019). This model is a binary classifier, and we use the ”$log\_softmax$” score of the final output layer as the detection score for AUROC calculation.
5. Experiment Results
Dataset APPS MBPP
Methods CodeLlama StarChat GPT-3.5 GPT-4 Avg. CodeLlama StarChat GPT-3.5 GPT-4 Avg.
GPTZero 52.71 56.25 53.68 58.24 55.22 59.53 60.82 57.29 61.55 59.80
OpenAI 56.32 50.08 48.48 55.81 52.67 48.81 47.40 43.31 46.44 46.49
Using CodeLlama as Detector LLM
log $p(x)$ 66.14 59.40 64.58 59.27 62.35 50.70 53.84 63.05 53.35 55.24
LogRank 69.79 61.54 67.31 62.89 65.38 60.76 58.56 68.05 58.91 61.57
Rank 52.17 48.63 50.77 48.04 49.90 25.99 35.75 42.03 36.33 35.03
Entropy 58.91 54.71 61.49 55.87 57.75 37.22 44.90 50.88 43.40 44.10
DetectGPT 61.28 57.71 62.06 53.41 59.85 56.28 53.18 66.56 63.84 59.96
LRR 67.15 62.16 67.82 60.06 64.30 53.25 56.32 64.29 54.62 57.12
NPR 65.49 60.08 66.53 58.62 62.68 54.37 55.10 68.85 64.96 60.82
Ours $m=2$ 80.78 72.91 73.12 68.19 73.75 77.90 68.88 76.36 75.02 74.54
Ours $m=4$ 85.42 76.53 77.70 74.29 78.49 82.91 71.50 79.83 77.71 77.99
Ours $m=8$ 87.77 78.13 80.23 74.51 80.16 86.21 75.70 83.58 81.75 81.81
Using StarChat as Detector LLM
log $p(x)$ 66.41 65.27 65.54 62.18 64.85 55.81 64.86 69.91 60.17 62.69
LogRank 66.95 65.81 66.74 64.25 65.93 58.69 65.31 69.56 59.55 63.28
Rank 53.85 48.24 50.37 49.77 50.56 37.24 44.48 47.97 46.47 44.04
Entropy 56.55 55.43 59.60 55.30 56.72 39.03 48.72 55.35 47.22 47.33
DetectGPT 60.92 58.23 61.52 58.62 61.26 54.41 55.74 66.49 65.02 60.42
LRR 66.55 68.91 68.45 65.88 67.45 56.80 66.74 69.87 60.16 63.39
NPR 64.47 63.60 66.43 65.00 64.88 54.20 60.00 70.43 66.17 62.70
Ours $m=2$ 81.93 77.23 72.46 72.89 76.13 79.68 73.79 79.28 69.24 75.50
Ours $m=4$ 85.51 79.24 74.58 77.35 79.17 80.61 76.44 81.05 74.67 78.19
Ours $m=8$ 87.24 81.35 76.28 77.84 80.68 83.67 79.00 83.17 78.04 80.97
Using GPT-3.5-Turbo as Detector LLM
Ours $m=2$ 77.84 81.67 79.02 79.04 79.39 66.21 77.29 83.05 83.45 77.50
Ours $m=4$ 78.21 82.22 82.12 78.69 80.31 67.00 78.87 85.39 82.23 78.37
Ours $m=8$ 78.47 82.48 83.25 80.87 81.27 67.66 79.23 86.23 84.00 79.28
Table 3. Main Results. The first two rows list AUROC scores of two detectors from GPTZero (2023) and OpenAI (2019) across two benchmarks and four Generation Tools. The following three sections detail the AUROC scores of our methods, with variations in $m$, along with seven other zero-shot detectors, using three Detector LLMs. CodeLlama is an abbreviation for CodeLlama-13B-Instruct, StarChat denotes StarChat-Alpha-15B, and GPT-3.5 refers to GPT-3.5-turbo. Red indicates the best performance, while blue indicates the best performance among all seven baselines.
5.1. Main Results
We list the AUROC score of all zero-shot detectors on the APPS and MBPP benchmarks in Table 3. Among the seven baselines, LogRank, LRR and NPR are the most compelling methods. However, the detection performance of baseline zero-shot detectors, including $\log p(x)$, DetecGPT, LRR and NPR drop significantly on the code benchmarks compared to their reported performance on detecting general texts (Mitchell et al., 2023; Su et al., 2023). This observation holds true regardless of the generation tools and detector LLMs used. The main reason lies in the fact that these SOTA zero-shot text detector are based on the log probability of tokens statistically. In the code domain, code exhibits a more uniform grammatical structure, often with a relatively fixed space for the next code tokens. As a result, there are many ”low-entropy” tokens in code, which significantly reduces the number of effective tokens that SOTA text detection methods can use to differentiate between human-written code and model-generated code, as illustrated in Figure 1. Furthermore, We observe that the previous baselines demonstrate improved performance when the Detector LLM and the Generation LLM are identical, and exhibit a decline in performance when they do not match.
As shown in Table 3, our proposed methods significantly outperform previous methods by all four prevalent code generation tools, with a 20.5% improvement on the APPS benchmark and 29.1% improvement on the MBPP benchmark. This improvement is primarily due to our proposed methods detecting synthetic code from a more holistic perspective instead of relying on token-wise scores. The holistic code rewriting and similarity measurement approach effectively avoids the inherent shortcomings of previous methods when applied to the code domain, while also leveraging the tendency of LLMs to select the most confident and likelihood code during generation.
It is worth noting that our method can achieve overwhelming detection performance with only $2$ rewriting for estimating the similarity expectation. Furthermore, using more rewriting ($m=4$ and $m=8$) can lead to better performance.
5.2. Effectiveness of Detector LLM
We observed that using GPT-3.5-turbo as the detector LLM for rewriting achieves optimal performance in detecting code generated by StarChat, GPT-3.5-turbo, and GPT-4, except for CodeLlama. Optimal performance in detecting code generated by CodeLlama is attained when using CodeLlama as the LLM detector.The primary reason for this is that among the four models (CodeLlama, StarChat, GPT-3.5-turbo and GPT-4), StarChat is fine-tuned on an instruction dataset distilled from GPT-3.5-turbo and GPT-4. Therefore, these three models may share a closer distribution. In contrast, CodeLlama is instruction-tuned on a dataset constructed by self-instructing Llama-2 (Touvron et al., 2023). We suspect that the suboptimal performance in detecting CodeLlama-generated code using GPT-3.5-turbo is due to the differences in the instruction datasets’ distributions.
Regarding the optimal use of GPT-3.5-turbo over StarChat for detecting StarChat-generated code, our observation of the rewritten code generated by both GPT-3.5-turbo and StarChat revealed that StarChat tends to oversimplify the rewritten code in complex problems, while GPT-3.5-turbo provides more accurate rewritten code.
5.3. Comparison to Supervised Detector
In Table 3, the first two rows represent two supervised detectors, GPTZero and OpenAI-Detector. We compare them with the zero-shot detectors and find that GPTZero and OpenAI-Detector fail to provide useful detection for code content, performing similarly to random guessing. Despite being trained on millions of labeled samples, these supervised detectors may only capture a small fraction of code content. This observation suggests that supervised detectors are prone to overfitting the training distribution and struggle to generalize to new domains without adaptive tuning. Furthermore, the proposed zero-shot detection methods demonstrate promising generalization capability when evaluated on different code distributions and outperform supervised detectors.
5.4. Ablation Study
We conducted an ablation experiment to analyze the contributions of two primary components in our design, i.e., Code Rewriting and Similarity Measurement. We considered two ablation setting: first, replacing the Code Rewriting with in-fill perturbation following DetectGPT (Mitchell et al., 2023) while retaining the Similarity Measurement (Sim); second, retaining Code Rewriting (CR) and replacing the similarity measurement between original and rewritten code with token-wise score difference (i.e, LogProb, LogRank, Rank and Entropy). The results are presented in Table 4.
Dataset APPS
Methods CodeLlama StarChat GPT-3.5 GPT-4
CR + Sim 87.77 76.96 80.23 74.51
CR + LogProb 57.29 46.70 50.58 55.13
CR + LogRank 75.28 64.76 75.85 67.69
CR + Rank 62.34 55.42 60.59 59.34
CR + Entropy 60.39 59.42 68.20 59.35
Perturb + Sim 74.55 71.55 74.23 53.43
Table 4. Ablation Study. All results are reported when using CodeLlama as the Detector LLM.
The ablation results indicate that the ”CR + Sim” method significantly enhances the detection performance. Replacing either Code Rewriting or Similarity Measurement leads to a noticeable decrease in performance, underscoring the importance of these two components.
6. Model Analysis
We conduct comprehensive experiments to assess the generalization of the proposed methods under diverse settings, including the impact of generation prompts, decoding strategy, programming languages, code correctness, code length, and revised scenarios. In addition, we also provide a sensitivity test for the choice of $m$ and the similarity models to understand our methods’ effectiveness better. For all experiments in this section, we use GPT-3.5-turbo as the generation tool.
6.1. Choice of Similarity Model
In the method section, we exploit self-supervised contrastive learning (Gao et al., 2021) to train a better code similarity model based on GraphCodeBERT (GCB-SimCSE). To investigate the impact of the similarity model, we experiment with three other variants of code similarity models: UniXcoder (Guo et al., 2022) trained by SimCSE (Unix-SimCSE), the original GraphCodeBERT model with average pooling (GCB-avg) and OpenAI’s text embedding service (Text-ada-002) (Greene et al., 2022). The results on the APPS benchmark (generated by GPT-3.5-turbo) are listed in Table 5. The UniXcoder trained with self-supervised contrastive learning (Unix-SimCSE) attains the highest performance, yet using the other three alternatives still surpasses the previous zero-shot detectors. The experimental results indicate that our method is not dependent on one specific similarity model. Moreover, as UniXcoder is an enhanced version of GraphCodeBERT, the superior performance of Unix-SimCSE over our original GCB-SimCSE implies that the detection performance can be further improved by utilizing a more strong and robust similarity model.
Method CodeLlama StarChat GPT-3.5
Unix-SimCSE 81.21 77.84 86.37
GCB-SimCSE 80.23 76.28 83.25
GCB-avg 72.79 70.95 72.22
Text-ada-002 70.18 69.29 76.48
Table 5. Test with different similarity models. Text-ada-002 is OpenAI’s text embedding service for text/code search or clustering. We use CodeLlama, StarChat and GPT-3.5-turbo as Detector LLMs.
6.2. Impact of Generation Prompts
It is well known that the output of large language model can be significantly affected by prompts. Consequently, the prompts used to generate synthetic code have a substantial impact on the generated output. To evaluate the impact of generation prompts, we altered the prompts used to generate synthetic code in our benchmark by removing the chain of thought prompts and using direct prompting, i.e., asking the model to generate a solution directly. The difference between these two prompts are illustrated in Figure 4 and the results under two different prompts are presented in Figure 5.
Figure 4. Prompt templates for synthetic code generation. We use chain-of-thought prompting for reporting main results in Table 3. We use direct prompting as part of model analysis for the impact of generation prompts.
Figure 5. Impact of generation prompts.
The results in Figure 5 indicate that our proposed methods are more robust than previous zero-shot detectors when the generation prompts are varied across both benchmarks.
6.3. Impact of Decoding Strategy
Adjusting the temperature parameter in large language models balances output diversity and accuracy, with lower temperatures yielding more deterministic and consistent results, while higher temperatures produce more varied and creative outputs. We conducted an experiment on APPS benchmark where we varied the generator temperature across the range of $[0.2,0.4,0.8]$ while keeping the rewriting temperature constant at $0.8$. The results are shown in Figure 6. This temperature range was chosen based on the findings from Codex paper (Chen et al., 2021), which identified $T=0.2$ and $T=0.8$ were optimal for pass@1 and pass@100 rates, respectively. The results in Figure 6 demonstrate that our method exhibits superior consistency across different temperatures.
Figure 6. Impact of Decoding Strategy.
6.4. Generalizability to Different Programming Languages
To verify the generalizability of our method to different programming languages, we construct an additional C++ benchmark using the Code Contest dataset (Li et al., 2022b). We generate synthetic code using the same four-generation tools. The detection performance on Code Contest C++ is presented in Table 6. Our proposed method still achieves notable improvement compared to other zero-shot baselines on C++ benchmarks. Additionally, we observe that synthetic C++ code is easier to detect for both our method and other baselines when compared to Python, suggesting a more significant distribution gap between synthetic and human-written C++ code.
Dataset Code Contest C++
Methods CodeLlama StarChat GPT-3.5 GPT-4
[…]
LogRank 59.35 57.12 66.68 62.43
Rank 54.67 52.19 60.29 54.21
Entropy 43.83 39.52 53.74 55.50
DetectGPT 62.99 60.45 75.47 64.05
Ours $m=8$ 89.87 83.42 90.82 88.85
Table 6. Detection Results on C++. All results are reported when using CodeLlama as the Detector LLM.
6.5. Detecting Revised Synthetic Code
In real-world scenarios, humans may revise synthetic code before using it. Naturally, this raises the question of whether zero-shot synthetic code detectors remain effective when a human user makes minor revisions to the generated code. We consider the most straightforward code modification: identifier renaming. This does not affect the functionality of the original code and even does not require the user to understand the code’s logic. To mimic identifier renaming, we extract all the identifier names and randomly replacing $10\%$, $20\%$, and $50\%$ fraction with ”$var\_i$”. Then we evaluate all the zero-shot detectors on the revised synthetic code to observe how performance changes with increasing fractions of replaced identifiers. The results are plotted in Figure 7.
Figure 7. Detecting Revised Synthetic Code.
From Figure 7, it’s evident that the performance of all detectors declines as the replacement fraction increase, degrading to random guess when $50\%$ of the identifiers are replaced. However, our method consistently outperforms all other detectors across all replacement fraction levels. This phenomenon can be attributed to the behavior of GPT-3.5-turbo and StarChat when rewriting the revised synthetic code. They tend to restore ”$var\_i$” to variable names that contain code semantics, resulting in a lower similarity between the rewritten code and the revised synthetic code compared to the similarity between the rewritten code and the original synthetic code. We consider detection in this or more complex adversarial scenarios as a focus for future work.
6.6. Impact of Code Correctness
We consider that the distribution of correct code solutions likely differ from incorrect ones since the correct model-generated code is much closer to human-written code and is more difficult to detect. To explore the impact of code correctness, we separately present the detection AUROC for both correct and incorrect codes on MBPP benchmark when using CodeLlama as Detector LLM in Figure 8.
Figure 8. Impact of code correctness.
From the results, we observe that both with our method and other baselines, detecting correct code is more challenging than detecting incorrect code. Nonetheless, our method consistently outperforms other baselines when it comes to correct code.
6.7. Impact of Code Length
We also investigate how the length of code snippet impacts the detection performance of the zero-shot synthetic code detectors. We divide the samples in the APPS benchmark into five groups according to the code string length. Each group has an equal number of synthetic and human-written code snippets. The results are plotted in Figure 9.
Figure 9. Impact of Code Length.
In general, the performance of all detectors improves as the code length falls within the 0-1000 character range. The $\log p(x)$, LogRank, and Entropy methods achieve optimal detection performance on code snippets between 600-1000, while exhibiting inferior performance on code lengths smaller than 200. DetectGPT’s performance surpasses ours when the code exceeds 1000 characters. However, DetectGPT is inferior to the other baselines for code length less than 600, the range where majority ($70\%$) of code snippets lie. Our method displays strong performance across all ranges and is more robust to changes in code length when compared to the other baselines.
6.8. Choice of $m$
In our primary experiment, we set the number of rewriting $m$ in the range of $[2,4,8]$ due to limited computational resources. However, increasing $m$ can reduce randomness in code sampling and enhance the accuracy of expectation estimation. To investigate this, we conduct experiments by setting the maximum value of $m$ to 32 on the APPS and MBPP benchmarks (generated by GPT-3.5-turbo) and plotted the detection AUROC against changes of $m$ in Figure 10. The result indicate that the detection performance increases with $m$ with slight fluctuation and saturate around 20 - 32 rewrites.
Figure 10. Choice of $m$. CodeLlama is used as Detector LLM.
7. Conclusion
In this paper, we identified the gap when applying synthetic content detectors designed for general texts to code content, and proposed a novel zero-shot synthetic code detector via code rewriting and similarity measurement. We utilize the similarity between rewritten and original code as an indicator for detecting synthetic code. Our proposed method addresses the unique challenges posed by the massive ”low-entropy” tokens in code content. Through exhaustive experiments, we uncovered several findings regarding synthetic code detection. The proposed method demonstrates superior performance and robustness in detecting synthetic code compared to other zero-shot baselines.